OpenStudy (k12awesomeness):

Medal and Fan!!! Which equation is represented by the table? x y (x, y) 0 3 (0, 3) 1 5 (1, 5) 2 7 (2, 7) A. y = x + 3 B. y = 5x C. y = 2x + 3 D. y = 3x + 2

OpenStudy (k12awesomeness):

@scoobydo @Hauntedwoodsgal @pink33

OpenStudy (texaschic101):

find the slope...only 1 answer choice has the correct slope. can u do that ?

OpenStudy (k12awesomeness):

I'm sorta just begging this stuff and i'm not very good at it, so no... I can't do that.

OpenStudy (scoobydo):

A or D

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (0,3)...x1 = 0 and y1 = 3 (1,5)...x2 = 1 and y2 = 5 now its just a matter of subbing...can u tell me the slope now ?

OpenStudy (k12awesomeness):

would it be D. ?

OpenStudy (texaschic101):

no..not d. u find the slope yet ?

OpenStudy (scoobydo):

me to

OpenStudy (scoobydo):

D is the right answer I think

OpenStudy (k12awesomeness):

I'm tryin to find the right slope but it's confusing me.

OpenStudy (texaschic101):

not d. just find the slope and we will work from there

OpenStudy (texaschic101):

slope = (5 - 3) / (1 - 0) slope = ?

OpenStudy (k12awesomeness):

Would the slope be 2 ?

OpenStudy (texaschic101):

yes...and in y = mx + b...slope is in the m position. so which answer has a 2 in the m position ?

OpenStudy (k12awesomeness):

C ?

OpenStudy (texaschic101):

correct..that is ur answer

OpenStudy (k12awesomeness):

thx

OpenStudy (texaschic101):

yw